# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v2

@@ Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects circuits, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory and Linear Recurrent Unit networks present more tendency to introduce audio artifacts, in particular, the Linear Recurrent Unit, which resulted in the least appropriate modeling techniques.
@@
The emulation of analog musical devices has become incredibly pervasive in digital commercial products. The distinct sound characteristics of both contemporary and old analog audio devices have captivated producers and musicians. Analog effect exists as standalone units as well as integrated with sound synthesizers. There are different types of popular effects such as filters, distortion, overdrive, compressor/limiter, as well as delay-based effects like chorus, flanger, phaser, and reverb. Analog circuits found in audio effects depend on intricately designed circuits that combine discrete passive and active electronic components to generate and modify electrical signals. Key components, such as operational amplifiers and diodes, are utilized in a variety of analog effects. These components present nonlinear behaviors that contribute to the unique sound of these devices. Accurate modeling of such nonlinearities is essential to replicate the behavior of analog audio devices. Consequently, the development of digital audio effects capable of simulating analog counterparts has become a crucial and valuable area of study within digital audio signal processing, often referred to as Virtual Analog (VA) modeling.
@@
This article presents a comparative study that applies recent advancements in machine learning to audio effects modeling. Machine learning has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Although hybrid solutions (referred to as gray-box approaches) have also been proposed, which offer their advantages [1], black-box approaches can still be appealing in some contexts. Gray-box approaches can incorporate physics knowledge [2] or measurements from systems [3]. While the former may not always be accurate enough, the latter may not be feasible for certain tasks. Universal black-box models have already been proposed [4], but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, needed to offer users runtime control over the effect’s parameters.
@@
In this study, we compare different recurrent artificial neural network architectures for modeling a selection of audio effects. In particular, we considered overdrive, saturator, low-pass filter, equalizer, and compressor. We selected audio effects with different complexities and time dependencies to investigate how these different architectural characteristics influence the results when learning directly from raw audio.
@@
For black-box modeling purposes, Recurrent Neural Networks (RNNs) are often preferred among the different types of artificial neural networks. RNNs utilize internal states, which function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially, when the inputs are limited to a small segment of digital audio samples, and the network does not have a large amount of input information to infer, as can happen with feed-forward architectures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution. This similarity to stateful machine learning architectures is often appreciated, as it allows for effective modeling of these systems.
@@
In this study, we do not consider Convolutional Neural Networks (CNNs) [10] although their integration with RNNs has the potential to enhance accuracy [11]. However, when working with small input vectors and considering long-term dependencies, an RNN-based structure with the ability to encode the signal’s history using the states is a particularly appealing feature. Therefore, in this study, our focus is solely on recurrent layers, and we do not include layers of different types from recurrent and fully connected layers.
@@ 2 Background
Physics-based modeling is used when comprehensive knowledge about the system is available. In the case of analog audio effects, this requires knowing the schematics of the circuits, and specific electronic components populating the system. In addition, the circuit and components need to be mathematically modeled and solvable. A numerical scheme has to be formulated and the obtained discrete equations have to be solved iteratively to produce the digital output sound. In the case of an implicit ordinary differential equations system, the Newton-Raphson (NR) iterative method is generally utilized to find the solutions. NR converges quadratically although this is not always guaranteed. As a consequence, physics-based models generally present a high computational cost.
@@
Common examples of data-driven approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in [13]. These approaches, for the reason of simplification, model linear and nonlinear behavior separately. Using this strategy two block types can be identified, linear dynamic blocks, and static nonlinearities.
Another data-driven approach for modeling analog circuits involves utilizing artificial neural networks (ANNs) [14]. ANNs are composed of interconnected units, where each connection can transmit a signal to other artificial neurons. These neurons receive and process signals, and they can transmit the signals to connected neurons. The signal transmitted through a connection is represented by a real number, and the output of each neuron is determined by a nonlinear function applied to the sum of its weighted inputs. Neural networks are trained by optimizing the network’s parameters to minimize the error between the predicted output and the actual target values within a given dataset. Gradient-based techniques, such as backpropagation, are commonly employed to estimate the network’s weights. Also in this scenario, it can be beneficial to incorporate a specific block that focuses on the ”unknown” behavior, while separately utilizing a different block to handle the ”known” aspects that have already been accurately modeled.
@@ 2.1 Virtual Analog Modeling
Similarly to a fully connected feedforward network, convolutional networks’ ability to track the past information in the signal depends on the receptive field, a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier [41], which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ( and ). Increasing the dilation factors increases in turn the receptive field and the ability of the networks when dealing with long dependencies while in turn not increasing the number of layers. In this case, the TCNs have been applied to kHz sampled audio files. A TCN block consists of a convolutional layer, batch normalization, conditioning using FiLM, and PRelu activation function. Authors compared networks with , , and TCN blocks, convolutional channel for each convolutional layer, and , , as the kernel size. The resulting models present receptive fields of , , and ms, respectively. The model with a receptive field of ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
@@
A model, that used a combination of different deep learning architectures, was investigated in [42]. In this case, the model is divided into three parts; an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. It also generates a residual connection which facilitates the reconstruction of the waveform by the back-end. The adaptive front-end consists of a convolutional encoder. It contains two CNN layers, one pooling layer, and one residual connection. The first layer has filters and channels and is followed by the absolute value as a nonlinear activation function. The second layer has filters of size . The latent-space DNN contains two layers. The first layer is based on locally connected dense layers of units and the second layer consists of a fully connected layer of units. Both of these layers are followed by the softplus function. The DNN modifies the latent representation of the input audio into a new latent representation which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. DNN-SAAF consists of four fully connected dense layers of 128, 64, 64, and 128 hidden units respectively. All dense layers are followed by the softplus function except the last layer. The SAAF [43] is used as the nonlinearity for the last layer. The model is first trained without the latent-space DNN and DNN-SAAF, and then these modules are incorporated for a second training. The loss function to be minimized is the mean absolute error (MAE). In both training procedures, the input and target audio are sliced into frames of samples with a hop size of samples. Distortion, overdrive, and EQ effects have been modeled using this method, with a sampling rate equal to kHz.
@@
In [4] the previously described architecture is further explored and compared with different variations and with the CNN architecture based on [38]. The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. The input frame size was sampled with a hop size of samples. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators [44]. Here, the loss function to be minimized is the sum of MAE in the time domain and MSE of the log power magnitude spectra of the target and output. These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, this architecture lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e. fixed parameters).
@@
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier [45], where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks consisted of or units, while the networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The number of past samples can differ from the input and output signal, and in this case, the experiments considered , , or input samples and or output samples. The work presents preliminary experiments and although the models resulted in poor performance, it demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs. Larger networks generated lower prediction errors and the overdrive case resulted in more challenges to model accurately.
@@
RNNs, in particular LSTMs, have been also compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals [46]. The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units (, , and ). The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model, while requiring significantly less computational cost and latency, since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when increasing the number of units. In addition, listening tests showed LSTM networks with units performing similarly to TCN with ms of latency [41].
@@
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in [47], which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar although the ML implementation has slightly damped high frequencies. The architecture consists of a GRU layer of units and one linear fully connected layer with one unit as the output layer. Also here a single variable control parameter of the audio effect is added as an extra input, to condition the inference of the networks.
@@
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor such as the TubeTech CL1B compressor [11], to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder the other as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, on processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks resulting in significantly better accuracy. The dataset, in this case, was at kHz sampling rate. The authors initially considered two control parameters: threshold and ratio, while attack and release parameters were fixed. In [50], the latter parameters were included. The encoder of the model was replaced with a convolutional layer and the model was compared to TCNs as well. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
@@
Other works consider hybrid approaches. In [3], fully connected neural networks exploit measurements within the circuits. Here, the structure includes to layers, embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, not always accessible, and is moving away from black-box models. Continuing towards gray box approaches, [51] is based on the Differentiable Digital Signal Processing (DDSP) framework, and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier, and a nonlinear guitar overdrive circuit in [52]. The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are in this way trainable and derived from the analysis of the circuit. This approach is also used in [53]. In this case, a cascade of biquad filters is trained employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit [54], showing higher accuracy and similar computation time when compared to traditional physics-based models. In [55], the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in [3]. Neural networks have been also investigated in the case of wave digital filter formulations [2]. A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in [56], authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is [57], where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
@@
Gray box methods have the advantage of helping the learning process with knowledge, potentially leading to more accurate and less computationally demanding models. On the other hand, internal measurement of circuit components is not always feasible because invasive or difficult to achieve. Infusing external knowledge could not be trivial. For example, in the case of an optical compressor, the attack and release time are governed by a lighting system whose behavior is difficult to model.

### v3

@@ Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects circuits, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory networks present more tendency to introduce audio artifacts.
@@
The emulation of analog musical devices has become incredibly pervasive in digital commercial products. The distinct sound characteristics of both contemporary and old analog audio devices have captivated producers and musicians. nalog effect exists as standalone units as well as integrated with sound synthesizers. There are different types of popular effects such as filters, distortion, overdrive, compressor/limiter, as well as delay-based effects like chorus, flanger, phaser, and reverb. Analog circuits found in audio effects depend on intricately designed circuits that combine discrete passive and active electronic components to generate and modify electrical signals. Key components, such as operational amplifiers and diodes, are utilized in a variety of analog effects. These components present nonlinear behaviors that contribute to the unique sound of these devices. Accurate modeling of such nonlinearities is essential to replicate the behavior of analog audio devices. Consequently, developing digital audio effects capable of simulating analog counterparts has become a crucial and valuable area of study within digital audio signal processing, often referred to as Virtual Analog (VA) modeling.
@@
This article presents a comparative study that applies recent advancements in machine learning to audio effects modeling. Machine learning has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Although hybrid solutions (referred to as gray-box approaches) have also been proposed, which offer their advantages [1], black-box approaches can still be appealing in some contexts. Gray-box approaches can incorporate physics knowledge [2] or measurements from systems [3]. While the former may not always be accurate enough, the latter may not be feasible for certain tasks. Universal black-box models have already been proposed [4], but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
@@
In this study, we compare different recurrent artificial neural network architectures for modeling a selection of audio effects. In particular, we considered overdrive, saturator, low-pass filter, equalizer, and compressor. We selected audio effects with different complexities and time dependencies to investigate how these architectural characteristics influence the results when learning directly from raw audio.
@@
For black-box modeling purposes, Recurrent Neural Networks (RNNs) are often preferred among the different types of artificial neural networks. RNNs utilize internal states, which function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially when the inputs are limited to a small segment of digital audio samples and the network does not have a large amount of input information to infer, as can happen with feed-forward architectures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution. This similarity to stateful machine learning architectures is often appreciated, as it allows for effective modeling of these systems.
@@
In this study, we consider Convolutional Neural Networks (CNNs) [10] only when integrated with RNNs [11]. However, when working with small input vectors and considering long-term temporal dependencies, an RNN-based structure with the ability to encode the signal’s history using the states is a particularly effective feature. Therefore, in this study, our focus is on models based on recurrent layers due to their effectiveness in capturing temporal dependencies, and we do not include convolutional-based networks, which are typically more suited for spatial pattern recognition.
@@ 2 Background
Physics-based modeling is used when comprehensive knowledge about the system is available. In the case of analog audio effects, this requires knowing the schematics of the circuits and specific electronic components populating the system. In addition, the circuit and components need to be mathematically modeled and solvable. A numerical scheme has to be formulated, and the obtained discrete equations have to be solved iteratively to produce the sound of the digital output. In the case of an implicit ordinary differential equations system, the Newton-Raphson (NR) iterative method is generally utilized to find the solutions. NR converges quadratically, although this is not always guaranteed. As a consequence, physics-based models generally present a high computational cost.
@@
Common examples of data-driven approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in [13]. These approaches, for the reason of simplification, model linear and nonlinear behavior separately. Using this strategy, two block types can be identified: linear dynamic blocks and static nonlinearities.
Another data-driven approach for modeling analog circuits involves utilizing artificial neural networks (ANNs) [14]. ANNs are composed of interconnected units, where each connection can transmit a signal to other artificial neurons. These neurons receive and process signals and can transmit them to connected neurons. A real number represents the signal transmitted through a connection, and the output of each neuron is determined by a nonlinear function applied to the sum of its weighted inputs. Neural networks are trained by optimizing the network’s parameters to minimize the error between the predicted output and the actual target values within a given dataset. Gradient-based techniques, such as backpropagation, are commonly employed to estimate the network’s weights. Also, in this scenario, it can be beneficial to incorporate a specific block that focuses on the ”unknown” behavior while separately utilizing a different block to handle the ”known” aspects that have already been accurately modeled.
@@ 2.1 Virtual Analog Modeling
Similarly to a fully connected feedforward network, convolutional networks’ ability to track the past information in the signal depends on the receptive field; a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier [41], which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ( and ). Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers. In this case, the TCNs have been applied to kHz sampled audio files. A TCN block consists of a convolutional layer, batch normalization, conditioning using FiLM, and PRelu activation function. Authors compared networks with , , and TCN blocks, convolutional channel for each convolutional layer, and , , as the kernel size. The resulting models present receptive fields of , , and ms, respectively. The model with a receptive field of ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
@@
A model, that used a combination of different deep learning architectures, was investigated in [42]. In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. It also generates a residual connection, which facilitates the reconstruction of the waveform by the back-end. The adaptive front-end consists of a convolutional encoder. It contains two CNN layers, one pooling layer, and one residual connection. The first layer has filters and channels and is followed by the absolute value as a nonlinear activation function. The second layer has filters of size . The latent-space DNN contains two layers. The first layer is based on locally connected dense layers of units and the second layer consists of a fully connected layer of units. Both of these layers are followed by the softplus function. The DNN modifies the latent representation of the input audio into a new latent representation which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. DNN-SAAF consists of four fully connected dense layers of 128, 64, 64, and 128 hidden units, respectively. All dense layers are followed by the softplus function except the last layer. The SAAF [43] is used as the nonlinearity for the last layer. The model is first trained without the latent-space DNN and DNN-SAAF, and then these modules are incorporated for a second training. The loss function to be minimized is the mean absolute error (MAE). In both training procedures, the input and target audio are sliced into frames of samples with a hop size of samples. Distortion, overdrive, and EQ effects have been modeled using this method, with a sampling rate equal to kHz.
@@
In [4] the previously described architecture is further explored and compared with different variations and with the CNN architecture based on [38]. The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. The input frame size was sampled with a hop size of samples. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators [44]. Here, the loss function to be minimized is the sum of MAE in the time domain and MSE of the log power magnitude spectra of the target and output. These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, this architecture lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
@@
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier [45], where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks consisted of or units, while the networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The number of past samples can differ from the input and output signal, and in this case, the experiments considered , , or input samples and or output samples. The work presents preliminary experiments and although the models resulted in poor performance, it demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs. Larger networks generated lower prediction errors, and the overdrive case resulted in more challenges in modeling accurately.
@@
RNNs, in particular LSTMs, have been also compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals [46]. The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units (, , and ). The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when increasing the number of units. In addition, listening tests showed LSTM networks with units performing similarly to TCN with ms of latency [41].
@@
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in [47], which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar although the ML implementation has slightly damped high frequencies. The architecture consists of a GRU layer of units and one linear fully connected layer with one unit as the output layer. Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
@@
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor [11], to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, on processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input, which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy. The dataset, in this case, was at kHz sampling rate. The authors initially considered two control parameters, threshold and ratio, while attack and release parameters were fixed. In [50], the latter parameters were included. The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
@@
Other works consider hybrid approaches. In [3], fully connected neural networks exploit measurements within the circuits. Here, the structure includes to layers embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, which are not always accessible, and is moving away from black-box models. Continuing towards gray box approaches, [51] is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in [52]. The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are, in this way, trainable and derived from the analysis of the circuit. This approach is also used in [53]. In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit [54], showing higher accuracy and similar computation time when compared to traditional physics-based models. In [55], the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in [3]. Neural networks have also been investigated in the case of wave digital filter formulations [2]. A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in [56], authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is [57], where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
@@
Gray box methods have the advantage of helping the learning process with knowledge, potentially leading to more accurate and less computationally demanding models. On the other hand, internal measurement of circuit components is not always feasible because it is invasive or difficult to achieve. Infusing external knowledge could not be trivial. For example, in the case of an optical compressor, the attack and release time are governed by a lighting system whose behavior is difficult to model.